# Krylov Complexity for Open Quantum System: Dissipation and Decoherence

### v1

@@ 1 Introduction
The paper is organized as follows. In Sec. (2), we outline the details of Krylov complexity and its computation from the moments of the two-point function. In Sec. (3), we discuss the damped harmonic oscillator model and present its Krylov complexity results. In Sec. (4), we introduce the Caldeira-Leggett master equation, its assumptions, and its features, and analyze the associated Krylov complexity and its implications. Finally, in Sec. (5), we summarize our findings and discuss their broader implications for future research. In the Appendix (A), we have provided some additional details.
@@ 2 Krylov Complexity
Krylov Complexity for Open Quantum System
For open quantum systems governed by a Lindbladian $\mathcal{L}$, which describes non-unitary evolution due to environmental coupling, the standard Lanczos algorithm for Hermitian operators is insufficient. Instead, the bi-Lanczos algorithm is required to handle the non-Hermitian nature of the Lindbladian Bhattacharya_2023 ; Chakrabarti:2025hsb . This algorithm constructs a bi-orthogonal Krylov basis, yielding a tridiagonal matrix representation with complex diagonal elements $a_{n}$ and off-diagonal elements $b_{n}$ (superdiagonal) and $c_{n}$ (subdiagonal). These off-diagonals are generally complex and unequal ($b_{n}\neq c_{n}^{*}$), reflecting the dissipative dynamics. In moment-based methods for computing Krylov complexity, the product $b_{n}c_{n}$ often appears, but its complex nature can complicate calculations, especially when taking the square root for complexity metrics. A similarity transformation can symmetrize the off-diagonals to $\sqrt{b_{n}c_{n}}$, simplifying moment computations while preserving the spectrum.
The Lindbladian’s tridiagonal form in the bi-Lanczos basis is Bhattacharya_2023 ; Chakrabarti:2025hsb ; NSSrivatsa:2023pby :
$\mathcal{L}=\begin{pmatrix}a_{0}&b_{0}&0&\cdots\\ c_{0}&a_{1}&b_{1}&\ddots\\ 0&c_{1}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$,
where $a_{n},b_{n},c_{n}\in\mathbb{C}$. The goal is to transform $\mathcal{L}$ such that the off-diagonals become equal, specifically $\sqrt{b_{n}c_{n}}$, to facilitate moment-based Krylov complexity calculations gaaf2016infinitebilanczosmethodnonlinear .
@@
Define a diagonal matrix $D=\text{diag}(d_{0},d_{1},\dots)$ with $d_{0}=1$ and ratios $d_{k+1}/d_{k}=\sqrt{c_{k}/b_{k}}$, choosing the principal branch of the square root (argument in $(-\pi/2,\pi/2]$) for consistency. The transformed Lindbladian is $\mathcal{L}^{\prime}=D^{-1}\mathcal{L}D$, which preserves eigenvalues and has the form Yamamoto1968 :
@@
$\mathcal{L}^{\prime}=\begin{pmatrix}a_{0}&\sqrt{b_{0}c_{0}}&0&\cdots\\ \sqrt{b_{0}c_{0}}&a_{1}&\sqrt{b_{1}c_{1}}&\ddots\\ 0&\sqrt{b_{1}c_{1}}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$.
@@
For complex values of $\sqrt{b_{n}c_{n}}$ gaaf2016infinitebilanczosmethodnonlinear , these terms can be written in modulus–argument form. In this representation, the equal off-diagonal elements share constant phases, which are absorbed by the transformation matrix $D$. This redefines the basis up to an overall phase, so that only the moduli of the coefficients need to be considered. Note that, in dissipative systems, such as the SYK model with Lindblad operators Bhattacharjee:2023uwx , $b_{n}=c_{n}$ may be real and positive, simplifying to $\sqrt{b_{n}c_{n}}=b_{n}$. For general cases, the transformation symmetrizes the off-diagonals, aiding moment computations for Krylov complexity while accounting for dissipative effects, where the imaginary parts of $a_{n}$ typically suppress complexity growth.
@@
In this transformed basis the Lanczos coefficients $\{b_{n}\}$ are replaced by $\sqrt{b_{n}c_{n}}$ and are related to the moments $\{\mu_{n}\}$ as NSSrivatsa:2023pby :
$b_{n}c_{n}$ $=\frac{\tau_{n+1}(t)\tau_{n-1}(t)}{\tau_{n}^{2}(t)},\quad n\geq 0.$ (12)
@@
In the following, we first consider a damped harmonic oscillator, followed by the well-known Caldeira–Leggett model as our representative open quantum systems. We then apply the moments method to compute and study the evolution of Krylov complexity.
@@ 3 Damped Harmonic Oscillator
$\frac{dA}{dt}=i[H,A]+\sum_{i}\left(L_{i}^{\dagger}AL_{i}-\frac{1}{2}L_{i}^{\dagger}L_{i}A-\frac{1}{2}AL_{i}^{\dagger}L_{i}\right)\,.$ (15)
The first term generates unitary evolution, while the remaining terms account for transitions due to interactions with the bath. The operators $L_{i}$ are jump operators, with $L_{i}^{\dagger}AL_{i}$ driving quantum transitions and the terms $\frac{1}{2}L_{i}^{\dagger}L_{i}A$ and $\frac{1}{2}AL_{i}^{\dagger}L_{i}$ correcting for cases where no jump occurs Preskill1998 .
@@
From the Fig. (1), we observe that at early times Krylov complexity exhibits rapid growth, followed by a decay toward a low saturation value. This behavior contrasts with the harmonic oscillator results presented in the Fig. (8) of the Appendix (A). Unlike the harmonic oscillator, which is a closed system, the damped oscillator loses information to the environment without any backflow. For the harmonic oscillator, the complexity displays periodic oscillations between zero and one, reflecting the evolution of the position operator into the momentum operator and vice versa, as dictated by the canonical commutation relations. These periodic oscillations are absent in the damped case.
In line with the conjecture made in Liu_2023 , dissipation leads to a lower overall complexity and a reduced saturation value. We therefore identify the oscillatory decay of Krylov complexity to a low saturation value as a hallmark of dissipation. Furthermore, the maximum complexity reached is significantly smaller than in the harmonic oscillator, as can be seen from the Fig. (1) and (8). This suggests that, in the overdamped regime, the operator explores only a limited portion of the system space before dissipating into the bath, while in the underdamped regime, it explores more of the system space (compared to the overdamped case, as evidenced by its higher saturation value) prior to this loss. Nevertheless, in both cases the spread into the bath occurs on the same timescale, consistent with the observed saturation. This interpretation is further supported by the probability amplitude plots as shown in Fig. (2): in the underdamped case, the slower decay indicates a less rapid loss of information to the bath.
4 Caldeira-Leggett Model
@@
In the Schrödinger picture, the dynamics of the reduced system are described by the Caldeira-Leggett master equation BRE02 :
$\frac{d}{dt}\rho_{s}(t)=-\frac{i}{\hbar}[H_{s},\rho_{s}(t)]-\frac{i\gamma}{\hbar}[x,\{p,\rho_{s}(t)\}]-\frac{2m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho_{s}(t)]],$ (25)
where $H_{s}=\frac{p^{2}}{2m}+\frac{m\omega^{2}x^{2}}{2}$. The first term describes coherent dynamics, the second term, proportional to the relaxation rate $\gamma$, drives dissipation by damping momentum expectation values, and the third term, proportional to temperature $T$, induces decoherence through thermal fluctuations Schlosshauer_2019 ; BRE02 . The Caldeira-Leggett master equation is not directly in Lindblad form, as its dissipator is not positive definite. In the high-temperature limit, this can be remedied by adding a minimally invasive term $\left(-\frac{\gamma}{8mk_{B}T}[p,[p,\rho_{s}(t)]]\right)$ Ramazanoglu_2009 ; BRE02 ; Schlosshauer_2019 . The Lindbladian form facilitates analysis of operator dynamics in the adjoint form:
$\frac{d}{dt}A(t)=i\left[H_{s}+\frac{\gamma}{2}(xp+px),A\right]+\gamma\left(L^{\dagger}AL-\frac{1}{2}L^{\dagger}LA-\frac{1}{2}AL^{\dagger}L\right),$ (26)
@@
where $A=\sqrt{\frac{4mk_{B}T}{\hbar^{2}}}$ and $B=\sqrt{\frac{1}{4mk_{B}T}}$. The approximate time evolution for an operator $O(t)$ is:
@@
which in terms of creation and annihilation basis can be written as
$\frac{d}{dt}O$ $=-\omega^{2}(xp+px)+(-2\gamma-2\gamma AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a).$ (29)
Here the proportionality factor is taken to be unity. Factorizing the time dependence of the creation and annihilation operators as $a(t)=f^{*}(t)a$ and $a^{\dagger}(t)=f(t)a^{\dagger}$, and solving the differential equation we obtain the following time evolved operator
@@
$C_{1}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}=\frac{(-1+e^{\beta\omega})^{2}}{3(e^{\beta\omega}+1)^{2}}\,$, (33)
@@
$C_{3}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}n^{2}e^{-\beta(n+1/2)\omega}=\frac{1}{3+3e^{\beta\omega}}\,.$ (35)
@@ 4.1 Decoherence Function
As a result of the commutation relation the basis states $|n\rangle$ are not effected by the coupled dynamics. This gives us the following system bath state:
@@
It represents correlations between states describing the system and the corresponding states of the reservoir. Due to the interaction, the reservoir carries information on the system state. These coherences/correlations are still present in the system’s reduced density matrix:
@@
The quantity $\Gamma_{nm}(t)$ characterizes the behaviour of the off diagonal terms in the reduced density matrix and is known as the decoherence function. For the Caldeira-Leggett model, the decoherence function is BRE02 :
$\Gamma(t)$ $=-\frac{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}{8\sigma_{0}^{2}}\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}},$ (43)
$\tilde{\Gamma}(t)$ $=\frac{8\sigma_{0}^{2}\Gamma(t)}{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}=\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}},$ (44)
@@
From Fig. (3), it is clear that the decoherence rate saturates, as described in Schlosshauer_2019 . This occurs when $x-x^{\prime}$ reaches the maximum coherence length of the oscillator environment. This length is the characteristic scale over which the phase coherences of the environment, and hence the bath correlations, remain significant. The saturation of the decoherence function implies that the off-diagonal elements of the density matrix vanish. However, this feature is basis-dependent, and as noted earlier in the section, the Krylov basis is not the appropriate one for capturing it. We therefore suspect that Krylov complexity may be insensitive to decoherence, a possibility we will explore in the next subsection.
4.2 Krylov Complexity
As outlined in Sec. (2), we use the moments method to compute the Krylov complexity for the Caldeira–Leggett model. To investigate the roles of dissipation and correlation, we examine the time evolution of the Krylov complexity, shown in Fig. (4).

### v2

@@ 1 Introduction
The paper is organized as follows. In Sec. (2), we outline the details of Krylov complexity and its computation from the moments of the two-point function. In Sec. (3), we discuss the damped harmonic oscillator model and present its Krylov complexity results. In Sec. (4), we introduce the Caldeira-Leggett master equation, its assumptions, and its features, and analyze the associated Krylov complexity and its implications. Finally, in Sec. (5), we summarize our findings and discuss their broader implications for future research. In the Appendix, we have provided some additional details.
@@ 2 Krylov Complexity
Figure 1: Schematic diagram of the spreading of the operator, as it evolves, in operator space.
Krylov Complexity for Open Quantum Systems
For open quantum systems governed by a Lindbladian $\mathcal{L}$ (a generalized Liouvillian), which describes non-unitary evolution due to environmental coupling, the standard Lanczos algorithm for Hermitian operators is insufficient. Instead, there are a variety of ways in which one can treat such quantum systems using Krylov methods. Among them are the Arnoldi algorithm, Closed Krylov basis and the bi-Lanczos algorithm. The closed Krylov basis Liu_2023 is an identical treatment to the generalization of the closed system dynamics for a non-Hermitian initial operator, as was discussed in the previous sub-section. Both the Arnoldi and bi-Lanczos algorithms attempt to generalize the Lanczos algorithm for open systems. The Arnoldi algorithm does this by constructing an orthonormal basis; however, it turns out that in this basis, the Lindbladian is not tri-diagonal, instead it is in an upper Hessenberg form. On the other hand the bi-Lanczos algorithm aims to preserve the tri-diagonal structure obtained in the Lanczos algorithm. This is done by evolving two initial vectors $|p_{0}\rangle$ and $|q_{0}\rangle$ by the adjoint Lindbladian $\mathcal{L^{\dagger}}$ and the Lindbladian $\mathcal{L}\,$, respectively. As such, it creates a bi-orthonormal basis instead of an orthonormal one,
$\mathrm{Kry}^{j}(\mathcal{L^{\dagger}},|p_{0}\rangle)$ $=\{|p_{0}\rangle,\mathcal{L^{\dagger}}|p_{0}\rangle,(\mathcal{L^{\dagger}})^{2}|p_{0}\rangle,\ldots\},$ (12)
$\mathrm{Kry}^{j}(\mathcal{L},|q_{0}\rangle)$ $=\{|q_{0}\rangle,\mathcal{L}|q_{0}\rangle,(\mathcal{L})^{2}|q_{0}\rangle,\ldots\},$ (13)
with the bi-orthonormality condition, $\langle q_{m}|p_{n}\rangle=\delta_{mn}$ Bhattacharya_2023; Chakrabarti:2025hsb. Unlike the Arnoldi algorithm, which leaves the Linbdladian in an upper Hessenberg form, the bi-Lanczos algorithm, as mentioned earlier, preserves the tri-diagonal structure with complex diagonal elements $a_{n}$ and off-diagonal elements $b_{n}$ (super-diagonal) and $c_{n}$ (sub-diagonal). Generated by equivalent statements of (12), by the action of $L$, $L^{{\dagger}}$ on the vectors.
$\mathcal{L^{{\dagger}}}|p_{j}\rangle$ $=b_{j}|p_{j-1}\rangle+a_{j}|p_{j}\rangle+c_{j+1}|p_{j+1}\rangle\,$, (14)
$\mathcal{L}|q_{j}\rangle$ $=c^{*}_{j}|q_{j-1}\rangle+a^{*}_{j}|q_{j}\rangle+b^{*}_{j+1}|q_{j+1}\rangle\,.$ (15)
Using this representation, the matrix elements of the Lindbladian can be written as $\langle q_{i}|\mathcal{L^{{\dagger}}}|p_{j}\rangle$, allowing it to be expressed as:
$\mathcal{L^{{\dagger}}}=\begin{pmatrix}a_{0}&b_{0}&0&\cdots\\ c_{0}&a_{1}&b_{1}&\ddots\\ 0&c_{1}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$ (16)
@@
where $a_{n},b_{n},c_{n}\in\mathbb{C}$. Using a similarity transformation, the off-diagonal entries can be written as a generalized Lanczos coefficient, $\tilde{b}_{n}=\sqrt{b_{n}c_{n}}$. Define a diagonal matrix $D=\text{diag}(d_{0},d_{1},\dots)$ with $d_{0}=1$ and ratios $d_{k}=\sqrt{c_{k}/b_{k}}$. The transformed Lindbladian is $\mathcal{L}^{\prime}=D^{-1}\mathcal{L}D$, which preserves eigenvalues and has the form Yamamoto1968:
@@
$\mathcal{L}^{\prime}=\begin{pmatrix}a_{0}&\sqrt{b_{0}c_{0}}&0&\cdots\\ \sqrt{b_{0}c_{0}}&a_{1}&\sqrt{b_{1}c_{1}}&\ddots\\ 0&\sqrt{b_{1}c_{1}}&a_{2}&\ddots\\ \vdots&\ddots&\ddots&\ddots\end{pmatrix}$. (17)
The benefit of this similarity transformation is that it allows us to generalize the moments method for closed systems to open systems. In general, the Lanczos coefficients $b_{n}$ are taken to be positive real numbers, with the diagonal coefficients generally being complex. For complex values of $\sqrt{b_{n}c_{n}}$ gaaf2016infinitebilanczosmethodnonlinear, these terms can be written in modulus–argument form. In this representation, the equal off-diagonal elements share constant phases, which are absorbed by the transformation matrix $D_{1}$.
$D_{1}=\mathrm{diag}\left(1,\exp\left\{i\phi_{1}\right\},\exp\left\{i(\phi_{1}+\phi_{2}\right\},\ldots,\exp\left\{i\sum_{n}\phi_{n}\right\}\right)$ (18)
Such that,
$\mathcal{L}^{\prime}=D^{-1}_{1}\mathcal{L}D_{1}\,.$ (19)
@@
This redefines the basis up to an overall phase, so that only the moduli of the coefficients need to be considered. This is possible as the diagonal coefficients $a_{n}$ are the ones that ensure the non-Hermitian nature of the Lindbladian Nandy_2025. In the absence of the complex diagonal coefficients, the Lindbladian/Liouvillian should be hermitian with the off-diagonal coefficients being, in the most general case, complex conjugate pairs caputa2021geometrykrylovcomplexity; Dymarsky_2021. Note that, in dissipative systems, such as the SYK model with Lindblad operators Bhattacharjee:2023uwx, $b_{n}=c_{n}$ may be real and positive, simplifying to $\sqrt{b_{n}c_{n}}=b_{n}$. For general cases, the transformation symmetrizes the off-diagonals, allowing for a form of a generalized moments method for Krylov complexity, which accounts for dissipative effects, where the imaginary parts of $a_{n}$ typically suppress complexity growth.
@@
As mentioned above, in this transformed basis, the Lanczos coefficients $\{b_{n}\}$ are replaced by $\tilde{b}_{n}=\sqrt{b_{n}c_{n}}$ and are related to the moments $\{\mu_{n}\}$ as NSSrivatsa:2023pby:
$\tilde{b}^{2}_{n}=b_{n}c_{n}$ $=\frac{\tau_{n+1}(t)\tau_{n-1}(t)}{\tau_{n}^{2}(t)},\quad n\geq 0\,.$ (20)
@@
In the following, we first consider a damped harmonic oscillator, followed by the well-known Caldeira–Leggett model as our representative open quantum systems. For both models, the Krylov basis is not large, and the $b_{n}$ Lanczos coefficients truncate fairly quickly, as is the case for the Harmonic Oscillator (shown in Appendix (A)). The Hilbert space of the system is infinite-dimensional, but because of the choice of the initial operator and associated algebra, the Lanczos $b_{n}$ coefficients terminate rapidly, as we will show later in Fig. (5). We then apply the moments method as described above to compute and study the evolution of Krylov complexity.
@@ 3 Damped Harmonic Oscillator
$\frac{dO(t)}{dt}=i[H,O(t)]+\sum_{i}\left(L_{i}^{\dagger}O(t)L_{i}-\frac{1}{2}L_{i}^{\dagger}L_{i}O(t)-\frac{1}{2}O(t)L_{i}^{\dagger}L_{i}\right)\,.$ (23)
The first term generates unitary evolution, while the remaining terms account for transitions due to interactions with the bath. The operators $L_{i}$ are jump operators, with $L_{i}^{\dagger}AL_{i}$ driving quantum transitions and the terms $\frac{1}{2}L_{i}^{\dagger}L_{i}O(t)$ and $\frac{1}{2}O(t)L_{i}^{\dagger}L_{i}$ correcting for cases where no jump occurs Preskill1998.
@@
From the Fig. (2), we observe that at early times Krylov complexity exhibits rapid growth, followed by a decay toward a low saturation value. This behavior contrasts with the harmonic oscillator results presented in the Fig. (10) of the Appendix (A). Unlike the harmonic oscillator, which is a closed system, the damped oscillator loses information to the environment without any backflow. This is captured in the presence of the complex $a_{n}$ Lanczos coefficients, as for the harmonic oscillator and the damped oscillator there is only one non-zero $b_{n}$ coefficient. For the harmonic oscillator, the complexity displays periodic oscillations between zero and one, reflecting the evolution of the position operator into the momentum operator and vice versa, as dictated by the canonical commutation relations. These periodic oscillations are absent in the damped case.
In line with the conjecture made in Liu_2023, dissipation leads to a lower overall complexity and a reduced saturation value. We therefore identify the oscillatory decay of Krylov complexity to a low saturation value as a hallmark of dissipation. Furthermore, the maximum complexity reached is significantly smaller than in the harmonic oscillator, as can be seen from the Fig. (2) and (10). This suggests that, in the overdamped regime, the operator explores only a limited portion of the system space before dissipating into the bath, while in the underdamped regime, it explores more of the system space (compared to the overdamped case, as evidenced by its higher saturation value) prior to this loss. Nevertheless, in both cases the spread into the bath occurs on the same timescale, consistent with the observed saturation. This interpretation is further supported by the probability amplitude plots as shown in Fig. (3). For the underdamped case, the slower decay indicates a less rapid loss of information to the bath.
4 Caldeira-Leggett master equation
@@
In the Schrödinger picture, the dynamics of the reduced system are described by the Caldeira-Leggett master equation:
$\frac{d}{dt}\rho_{s}(t)=-\frac{i}{\hbar}[H_{s},\rho_{s}(t)]-\frac{i\gamma}{\hbar}[x,\{p,\rho_{s}(t)\}]-\frac{2m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho_{s}(t)]]$ (33)
where once again $H_{s}$ is the standard Hamiltonian of the quantum harmonic oscillator. As seen from the derivation the terms in the master equation come directly from the dissipative and noise kernels respectively, as such the first term describes coherent dynamics, the second term, proportional to the relaxation rate $\gamma$, drives dissipation by damping momentum expectation values, and the third term, proportional to temperature $T$, induces decoherence through thermal fluctuations Schlosshauer_2019; BRE02. The Caldeira-Leggett master equation is not directly in Lindblad form, as its dissipator is not positive definite. In the high-temperature limit, this can be remedied by adding a minimally invasive term, as was seen in the discussion in the derivation of the C-L master equation Ramazanoglu_2009; BRE02; Schlosshauer_2019. The Lindbladian form facilitates analysis of operator dynamics in the adjoint form:
$\frac{d}{dt}O(t)=i\left[H_{s}+\frac{\gamma}{2}(xp+px),O(t)\right]+\gamma\left(L^{\dagger}O(t)L-\frac{1}{2}L^{\dagger}LO(t)-\frac{1}{2}O(t)L^{\dagger}L\right),$ (34)
@@
where the coefficients $A$ and $B$ are the following:
$A=\sqrt{\frac{4mk_{B}T}{\hbar^{2}}}\,,\quad B=\sqrt{\frac{1}{4mk_{B}T}}\,.$ (36)
In a similar manner to how we computed the operator evolution in the case of the damped oscillator. We begin with an initial operator $O(0)=p^{2}$, then the commutator structure of the RHS of the master equation (34) yields the following operator expression  [^1] The proportionality sign is there to indicate that the RHS of this expression does not have any time-dependent contribution.:
@@
Gathering terms and writing the identity term in terms of creation and annihilation operators, the expression can be written as
$\frac{d}{dt}O(t)$ $\propto-\omega^{2}(xp+px)+(-2\gamma-2\gamma AB)p^{2}+\gamma A^{2}(aa^{\dagger}-a^{\dagger}a)\,.$ (38)
Details of the derivation are given in Appendix (B). Solving the differential equation and gathering the terms together, we obtain the following time-evolved operator
@@
$C_{1}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}e^{-\beta(n+1/2)\omega}=\frac{(-1+e^{\beta\omega})^{2}}{3(e^{\beta\omega}+1)^{2}}\,$,
@@
$C_{3}$ $=\frac{1}{\mathrm{Tr}\{\rho_{\beta}O^{2}\}}\sum_{n}n^{2}e^{-\beta(n+1/2)\omega}=\frac{1}{3+3e^{\beta\omega}}\,.$
@@ 4.1 Decoherence Function
As a result of the commutation relation, the basis states $|n\rangle$ are not affected by the coupled dynamics. This gives us the following system bath state:
@@
It represents correlations between states describing the system and the corresponding states of the reservoir. Due to the interaction, the reservoir carries information on the system state. These coherence/correlations are still present in the system’s reduced density matrix:
@@
The quantity $\Gamma_{nm}(t)$ characterizes the behaviour of the off-diagonal terms in the reduced density matrix and is known as the decoherence function. For the Caldeira-Leggett model, the decoherence function is BRE02:
$\Gamma(t)$ $=-\frac{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}{8\sigma_{0}^{2}}\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}}\,$,
$\tilde{\Gamma}(t)$ $=\frac{8\sigma_{0}^{2}\Gamma(t)}{\Delta x^{2}+4\sigma_{0}^{4}\Delta p^{2}}=\frac{\langle I^{2}(t)\rangle}{\langle I^{2}(t)\rangle+\sigma_{0}^{2}G_{1}^{2}(t)+\frac{G_{2}^{2}(t)}{4m^{2}\sigma_{0}^{2}}}$ (50)
@@
From Fig. (4), it is clear that the decoherence rate saturates, as described in Schlosshauer_2019. This occurs when the separation between two system positions $x$ and $x^{\prime}$ reaches the maximum coherence length of the environment [^2] When the separation between $x$ and $x^{\prime}$ exceeds the coherence length, as the environment is performing a type of “measurement” on the system, it cannot distinguish any further between the points causing decoherence to saturate. The saturation of the decoherence function implies that the off-diagonal elements of the density matrix vanish. However, this feature is basis-dependent, and as noted earlier in the section, the Krylov basis is not the appropriate one for capturing it. We therefore suspect that Krylov complexity may be insensitive to decoherence, a possibility we will explore in the next subsection.
4.2 Lanczos Coefficients and Krylov Complexity
Now we will turn our attention to the Krylov complexity. Before presenting the results for it, we will first examine the nature of the Lanczos coefficients, which subsequently enter into the computation of the complexity. Note that the dimension of the Krylov space largely depends on the underlying algebra and the choice of the initial operator. For the cases considered in this paper, we start with a simple operator to probe the effects of dissipation and decoherence in these models. Below is a plot of the Lanczos coefficients $\tilde{b}_{n}$ vs n.n\,.
Figure 5: $\tilde{b}_{n}$ Lanczos coefficients for the C-L model, with $\omega=1$ and $\gamma=1/10$.
Similar to the results for the closed system (the quantum harmonic oscillator), and as expected, the Lanczos coefficients terminate quickly as a result of the algebra and initial operator choice. While the $a_{n}$ coefficients are complex and account for the dissipative features in the Krylov complexity, the $\tilde{b}_{n}^{\prime}s$ remain real. It is clear that, simply by examining the behavior of the Lanczos coefficients, we cannot obtain any significant information about the dynamics, especially regarding the role of dissipation. Hence we turn our attention to the Krylov complexity. As outlined in Sec. (2), we use the moments method to compute the Krylov complexity for the Caldeira–Leggett model. To investigate the roles of dissipation and correlation, we examine the time evolution of the Krylov complexity, shown in Fig. (6).